OpenStudy (anonymous):

if 18 is subtracted from twice the square of an integer, the result is equal to nine times the integer. the integer must be

OpenStudy (anonymous):

Okay for this type of problems. let us assume the integer number be x

OpenStudy (anonymous):

then square of number becomes x^2 and twice the square of number becomes 2x^2

OpenStudy (anonymous):

Let's give this integer a name. Let's call it x. What is happening to x here? You are subtracting 18 from twice the square of 'x', and that's equal to 9 times x? Can you set up an equation for that? @shw18

OpenStudy (anonymous):

now as he said that if 18 subtracted then, what we would have as equation is 2x^2 -18 right?

OpenStudy (anonymous):

Now that is equals to the nine times integer which 9x therefore 2x^2 -18 = 9x

OpenStudy (anonymous):

Solve for x and you will get two posible numbers and they are the solution

OpenStudy (anonymous):

2x ^{2}-18=9x ? @genius12

OpenStudy (anonymous):

Yes @shw18

OpenStudy (anonymous):

et x = the integer: 2x^2 - 18 = 9x 2x^2 - 9x - 18 = 0 Now either factor that or use the quadratic formula if you're lazy like me. x = -3/2 and 6, so x must be 6 because it has to be an integer. Done!

OpenStudy (anonymous):

after i hav my equation though where do i go from there